Solid sphere is rolling on the surface and velocity of c.o.m is V m/s get strike to another sphere of the same mass kept at rest. Nature of collision is elastic . Find out velocity of both the sphere .Just after the collision and also angular velocity. ....

Solid  sphere  is rolling  on  the  surface and  velocity  of  c.o.m is V m/s get strike  to  another  sphere  of  the  same mass kept at rest. Nature of collision  is elastic . Find out velocity  of  both  the  sphere  .Just  after  the collision  and  also  angular  velocity. .... 

Grade:12th pass

1 Answers

Vikas TU
5 years ago
Dear Student,
Since external force is absent here momentum will be conserved.
Let m be their masses. Then,
mv + 0 = mv1 + mv2
v = v1 + v2..............(1)
 
Since the collision is elastic, e = 1
1 = (v2 – v1)/v
or
v = v2 – v1.....................(2)
Solving both the equations, we get
2*v = 2*v2
v2 = v
and v1 = 0
Hence, the solid shpehere will keep rolling with its inital speed v and second sphere will be at rest again even after collision.